OpenStudy (anonymous):

BDFH is a rectangle with a width of 5 and a length of 12. One diagonal is 2x + 3y and the other diagonal is x + 8y. Find the values of x and y. x = 4 and y = 1 x = 4 and y = -1 x = 5 and y = 1 x = 5 and y = -1

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@uri

OpenStudy (raden):

hint : diagonal = sqrt(l^2 + w^2) l=length, w=width

OpenStudy (anonymous):

so wt do i do 5^2 + 12^3?

OpenStudy (raden):

5^2 = 5 * 5 = 25 and 12^2 = 12 * 12 = 144 add of them, and take squareroot after that

OpenStudy (anonymous):

so add 144 + 25 = 169

OpenStudy (raden):

sqrt(169) = ... ?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

bt thats not an answer option

OpenStudy (anonymous):

so wt do i do now

OpenStudy (raden):

yes, it just a start

OpenStudy (raden):

now, the diagonal represnted by equations : d=2x + 3y and d=x+8y and we knowed that the length of diagonals of rectangle is same, so 2x+3y=13 and x+8y=13 solve for x and y by elimination or subt method

OpenStudy (raden):

or u can do trial error from ur options

OpenStudy (anonymous):

i got it x= 5 y = 1

OpenStudy (raden):

yep, that's correct

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome